# Upsample or Upweight? Balanced Training on Heavily Imbalanced Datasets

### v2

@@ Abstract
Data availability across domains often follows a long-tail distribution: a few domains have abundant data, while most face data scarcity. This imbalance poses challenges in training language models uniformly across all domains. In our study, we focus on multilingual settings, where data sizes vary significantly between high- and low-resource languages. Common strategies to address this include upsampling low-resource languages (Temperature Sampling) or upweighting their loss (Scalarization). Although often considered equivalent, this assumption has not been proven, which motivates our study. Through both theoretical and empirical analysis, we identify the conditions under which these approaches are equivalent and when they diverge. Specifically, we demonstrate that these two methods are equivalent under full gradient descent, but this equivalence breaks down with stochastic gradient descent. Empirically, we observe that Temperature Sampling converges more quickly but is prone to overfitting. We argue that this faster convergence is likely due to the lower variance in gradient estimations, as shown theoretically. Based on these insights, we propose Cooldown, a strategy that reduces sampling temperature during training, accelerating convergence without overfitting to low-resource languages. Our method is competitive with existing data re-weighting and offers computational efficiency.
@@ 1 Introduction
The internet is full of information, ranging from common knowledge like famous landmarks and popular recipes to rare details such as local folklore and specialized scientific theories. Some data are widely available, while others are much harder to collect (Kandpal et al., 2023). Data availability across different domains is often long-tailed (Feldman, 2020; Feldman and Zhang, 2020), with few domains having an abundance of data and many rare domains where data is scarce. However, the standard language model training objective weighs each training instance equally, paying little attention to domains that suffer from data scarcity. The heavy mismatch in dataset sizes creates substantial challenges in training language models to be competent in all domains.
@@
Specifically, we focus on language modeling on a natural divide of domains with a heavy mismatch: different languages in multilingual language modeling. As a result of this mismatch, language models are often trained on corpora with an overwhelming amount of English and other high-resource languages (HRLs) and very few amounts of data for low-resource languages (LRLs) (Koehn and Knowles, 2017; Conneau et al., 2020; Xue et al., 2021). For example, the multilingual C4 corpus (Xue et al., 2021) contains 2733 billion English tokens but only 1 billion Swahili tokens. Uniformly sampling from the large, combined dataset would result in the language model optimized heavily towards performance on HRLs (e.g. English), sacrificing performance on LRLs (e.g. Swahili).
Two methods are often employed to address domain mismatches: Scalarization and Temperature Sampling. Scalarization adjusts the losses for individual domains by re-weighting them under uniform sampling (Zhou et al., 2021; Choi et al., 2023). In this case, we assign a larger weight to LRLs to emphasize their importance. Temperature Sampling, weights each training instance uniformly and handles the mismatch by over-sampling HRLs and/or down-sampling LRLs. (Aharoni et al., 2019; Wang et al., 2020b; Chung et al., 2023; Xue et al., 2021). Intuitively, Scalarization modifies the loss while Temperature Sampling modifies the dataset. Scalarization and Temperature Sampling are widely regarded as equivalent. Choi et al. (2023) denotes “we follow convention and implement Scalarization via proportional sampling”. Xie et al. (2023) and Fan et al. (2024) implement sampling probabilities by multiplying losses with per-domain re-normalized weights. The underlying assumption is that Temperature Sampling and Scalarization are equivalent, and we can use them interchangeably. However, to the best of our knowledge, this assumed equivalence has neither been proven nor disproven.
We closely investigate this assumed equivalency theoretically (§3.1). Specifically, we prove that although they are equivalent in full gradient descent (Theorem 1), Temperature Sampling induces lower variance in the context of stochastic gradient descent (Theorem 2). Moreover, the variance induced by scalarization increases as the approximated temperature increases or the domain distribution’s skewness increases (Theorem 3). Based on our theoretical results and connecting to the literature on lower variance between stochastic gradients accelerates convergence Sutskever et al. (2013), we make the following hypothesis:
@@
Temperature Sampling (TS) converges much faster than Scalarization (S) at higher temperatures or on heavily imbalanced domain distributions.
We empirically verify our hypothesis (§3.2) and find that Temperature Sampling does converge faster but is more prone to overfitting. We identify in TS, the temperature controls the speed of convergence and hence can be used as a control knob to adjust the convergence speed. We thus propose Cooldown: to use a large temperature initially for fast convergence, then decrease the temperature to prevent overfitting to the LRLs. Figure 1 illustrates the effectiveness of Cooldown, which significantly accelerates convergence on the LRL due to a high temperature (aggressive upsampling of LRLs) at the beginning of training and reduces overfitting to the LRL due to the lowering the temperature during training.
@@
  • Motivated by our findings, we propose Cooldown, a method to adjust the sampling temperature during training on unbalanced datasets. We show the effectiveness of Cooldown on two tasks: multilingual machine translation and multilingual language modeling (§4).
@@ 2.1 Notations and Task Description
We consider a model is trained on a collection of data $\mathcal{D}=\{x\}_{i=1}^{N}$ from $K$ domains $\mathcal{D}=\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup...\cup\mathcal{D}_{K}$. Here, “domain" refers to sources (Books, Wikipedia, code) for general language modeling or different languages (English, French, Swahili) in multilingual language modeling. The total training loss $\mathcal{J}(D)$ is the sum of the losses of each example $\mathcal{L}(x)$.
$\mathcal{J}(D)=\sum_{x\in\mathcal{D}}\mathcal{L}(x).$
@@
Naive aggregation often results in imbalanced performance across domains when high-resource domains dominate the aggregated loss. Scalarization solves this issue by assigning weights $\{w_{i}\}_{i=1}^{K}$ to each domain and aggregates the weighted sum of individual losses:
$\mathcal{L}_{\textrm{S}}=\sum_{x\in\mathcal{D}}w_{f(x)}\mathcal{L}(x),$
@@
Instead of assigning weights to losses, we can also sample more frequently from the low-resource domain to achieve balanced training. Temperature sampling achieves this by adjusting the probabilities of selecting instances from different domains based on their sizes. The sampling probability vector of each domain $p$ is given by:
$\forall i\in\{1,2,...,K\}:\;p(\mathcal{D}_{i};\tau)=\frac{|\mathcal{D}_{i}|^{\frac{1}{\tau}}}{\sum_{j=1}^{K}|\mathcal{D}_{j}|^{\frac{1}{\tau}}},$
@@
$\mathcal{L}_{\textrm{TS}}=\mathop{\mathbb{E}}_{\mathcal{D}_{i}\sim p}\left[\sum_{x\in\mathcal{D}_{i}}\mathcal{L}(x)\right].$
The common understanding is that Temperature Sampling is equivalent to Scalarization (Choi et al., 2023; Kurin et al., 2022; Xin et al., 2022), and we can use them interchangeably (Xie et al., 2023). In the next subsection, we will formalize this statement and show that these two are mathematically equivalent in full gradient descent. We will then show that they are not equivalent under stochastic gradient descent.
@@ 3.1 Theoretical Analysis
For any sampling temperature $\tau$, there exists a set of weights $\textbf{w}=\{w_{1},w_{2},...,w_{K}\}$ for the Scalarization loss such that this loss is equivalent to the Temperature Sampling loss, both computed based on the whole data $\mathcal{D}$.
@@
$w_{i}=\frac{\sum_{j=1}^{K}|D_{j}|}{|D_{i}|}\cdot\frac{|\mathcal{D}_{i}|^{\frac{1}{\tau}}}{\sum_{j=1}^{K}|\mathcal{D}_{j}|^{\frac{1}{\tau}}}=\frac{p(\mathcal{D}_{i};\tau)}{p(\mathcal{D}_{i};1)},$
then
$\mathcal{L}_{\textrm{S}}$ $=\mathop{\mathbb{E}_{x\sim D}}\left[w_{f(x)}\mathcal{L}(x)\right]$
$=\sum_{x\in\mathcal{D}}\frac{p(\mathcal{D}_{f(x)};1)}{|\mathcal{D}_{f(x)}|}w_{f(x)}\mathcal{L}(x)$
$=\sum_{i=1}^{K}p(\mathcal{D}_{i};\tau)\sum_{x\in\mathcal{D}_{i}}\frac{1}{|\mathcal{D}_{i}|}\mathcal{L}(x)$
$=\mathop{\mathbb{E}}_{\mathcal{D}_{i}\sim p(\cdot;\tau)}\left[\sum_{x\in\mathcal{D}_{i}}\frac{1}{|\mathcal{D}_{i}|}\mathcal{L}(x)\right]$
$=\mathbb{E}_{\mathcal{D}_{i}\sim p(\cdot;\tau)}[\mathbb{E}_{x\sim\mathcal{D}_{i}}[\mathcal{L}(x)]]=\mathcal{L}_{\textrm{TS}}.$
@@
Denote the total gradient as $\nabla\mathcal{L}_{S}=\nabla\mathcal{L}_{TS}$, and $\nabla\mathcal{L}(x)$ is the gradient with respect to a single datapoint $x$. We denote the stochastic gradient under Scalarization as $\nabla\mathcal{L}_{\textrm{S}}(x)=\{\nabla w_{f(x)}\mathcal{L}(x)|x\sim\mathcal{D}\}$, which means the stochastic gradient $\nabla w_{f(x)}\mathcal{L}(x)$ when $x$ is uniformly drawn from the dataset $\mathcal{D}$. Similarly, we denote the stochastic gradient under Temperature Sampling as $\nabla\mathcal{L}_{\textrm{TS}}(x)=\{\nabla\mathcal{L}(x)|x\sim\mathcal{D}_{i},\mathcal{D}_{i}\sim p(\mathcal{D}_{i};\tau)\}$. Then both $\nabla\mathcal{L}_{\textrm{S}}(x)$ and $\nabla\mathcal{L}_{\textrm{TS}}(x)$ are unbiased estimates of the total gradient.
@@
By Theorem 1, $\nabla\mathcal{L}_{s}=\mathop{\mathbb{E}}_{x\sim\mathcal{D}}\left[w_{f(x)}\nabla\mathcal{L}(x)\right]$, and $\nabla\mathcal{L}_{TS}=\mathbb{E}_{\mathcal{D}_{i}\sim p(\cdot;\tau)}[\mathbb{E}_{x\sim\mathcal{D}_{i}}[\nabla\mathcal{L}(x)]]$.
We conclude that
$\mathop{\mathbb{E}_{\mathcal{D}_{i}\sim p(\mathcal{D}_{i};\tau)}\left[\mathbb{E}_{x\sim\mathcal{D}_{i}}\left[\nabla\mathcal{L}(x)\right]\right]}$
$=\mathbb{E}_{x\sim\mathcal{D}}$ $[w_{f(x)}\nabla\mathcal{L}(x)].$
@@
Using the same notation in Corollary 1.1, we have $\mathrm{Var}(\nabla\mathcal{L}_{\textrm{S}}(x))\geq\mathrm{Var}(\nabla\mathcal{L}_{\textrm{TS}}(x))$.
We provide a detailed proof in Appendix B.
@@
The difference $\Delta=\mathrm{Var}(\nabla\mathcal{L}_{\textrm{S}}(x))-\mathrm{Var}(\nabla\mathcal{L}_{\textrm{TS}}(x))$ monotonically increases when $\tau\geq 1$.
@@
Figure 2: Variance of Scalarization $\sum_{i}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}$ by sampling temperature $\tau$. A large temperature or a skewed distribution of $\mathcal{D}$ induces a much larger variance for Scalarization. Distributions $\mathcal{D}_{i}\propto\frac{1}{i^{\alpha}}$. See Appendix C for details of the experiment setup.
@@ 3.2 Empirical Evidence
Figure 3 shows the variance between mini-batch gradients for Temperature Sampling and Scalarization. Figure 4 shows the validation loss curves by training iterations.
Figure 3: Gradient Norm distributions on En-{Cs, Ro} for sampling and Scalarization. Empirical results show that Scalarization induces a larger variance (2.25 >>> 0.62) between mini-batch gradients compared to Temperature Sampling.
@@
We found that temperature sampling converges faster than scalarization at all temperatures we experimented with. This supports our theory that Temperature Sampling lowers the variance in gradient estimation (Theorem 2), thus speeding up convergence. The larger the temperature, the larger the gap between the convergence speed of Temperature Sampling v.s. Scalarization. We conclude that when there is a heavy mismatch in sizes, it is better to use Temperature Sampling with a large temperature, confirming the intuition that when the LRL has very little data, we want to upsample it more aggressively. We summarize our findings below:

### v3

@@ Abstract
Data abundance across different domains exhibits a long-tailed distribution: few domains have abundant data, while most face data scarcity. Our work focuses on a multilingual setting, where available data is heavily skewed toward high-resource languages, creating significant imbalances in training data sizes across languages. This disparity challenges training language models to perform uniformly well in all languages. Two common strategies to address this issue are upsampling low-resource languages (Temperature Sampling) and upweighting their loss functions (Scalarization). These methods are often assumed to be equivalent, but this equivalence has not been rigorously established, prompting our investigation.
Through theoretical and empirical analysis, we identify when these two methods are equivalent and when they diverge. We prove that they are equivalent under full gradient descent but differ under stochastic gradient descent due to differences in gradient variance. Specifically, Temperature Sampling exhibits lower variance in gradient estimation, leading to faster convergence but a higher risk of overfitting. Based on these insights, we propose Cooldown, a strategy that starts by heavily upsampling low-resource languages to accelerate convergence and gradually reduces the upsampling to prevent overfitting—achieving the best of both worlds. Our method competes effectively with existing data re-weighting techniques while offering computational efficiency.
@@ 1 Introduction
Information on the internet ranges from common knowledge, such as famous landmarks, to rare details, such as local folklore and specialized scientific theories. Data availability across different domains is often long-tailed (Feldman, 2020; Feldman and Zhang, 2020; Kandpal et al., 2023), where very few domains have abundant data. However, the standard language model training objective weighs each training instance equally, paying little attention to domains that suffer from data scarcity. The heavy mismatch in dataset sizes creates substantial challenges in training language models to be competent in all domains.
@@
This work focuses on language modeling on a natural divide of domains with a heavy mismatch: different languages in multilingual language modeling. Multilingual language models are often trained on corpora with an overwhelming amount of English and other high-resource languages (HRLs) and tiny amounts of data for low-resource languages (LRLs) (Koehn and Knowles, 2017; Conneau et al., 2020; Xue et al., 2021). For example, the multilingual C4 corpus (Xue et al., 2021) contains 2733 billion English tokens but only 1 billion Swahili tokens. Uniformly sampling from the combined dataset would result in the language model optimized heavily towards performance on HRLs (e.g. English), sacrificing performance on LRLs (e.g. Swahili).
Two methods are often employed to address domain mismatches: Scalarization and Temperature Sampling. Scalarization adjusts the losses for individual domains by re-weighting them under uniform sampling (Zhou et al., 2021; Choi et al., 2023). In this case, we assign a larger weight to LRLs to emphasize their importance. Temperature Sampling, weights each training instance uniformly and handles the mismatch by over-sampling LRLs and/or down-sampling HRLs (Aharoni et al., 2019; Wang et al., 2020b; Chung et al., 2023; Xue et al., 2021). Intuitively, Scalarization modifies the loss while Temperature Sampling modifies the dataset. Scalarization and Temperature Sampling are widely regarded as equivalent. Choi et al. (2023) denotes “we follow convention and implement Scalarization via proportional sampling”. Xie et al. (2023) and Fan et al. (2024) implement sampling probabilities by multiplying losses with per-domain re-normalized weights. The underlying assumption is that Temperature Sampling and Scalarization are equivalent, and we can use them interchangeably. However, to the best of our knowledge, this equivalence has not been rigorously established.
We closely investigate this assumed equivalency in theory (§3.1). Specifically, we prove that although they are equivalent in full gradient descent (Theorem 1), Temperature Sampling induces lower variance in the context of stochastic gradient descent (Theorem 2). Moreover, the variance induced by scalarization increases as the approximated temperature increases or the domain distribution’s skewness increases (Theorem 3). Based on our theoretical results and connecting to the literature on lower variance between stochastic gradients accelerates convergence Sutskever et al. (2013); mccandlish2018empiricalmodellargebatchtraining, we make the following hypothesis:
@@
Temperature Sampling converges much faster than Scalarization at higher temperatures or on heavily imbalanced domain distributions.
We empirically verify our hypothesis (§3.2) and find that Temperature Sampling does converge faster but is more prone to overfitting. We identify that the temperature controls the speed of convergence and hence can be used as a control knob to adjust the convergence speed. We thus propose Cooldown: to use a large temperature initially for fast convergence, then decrease the temperature to prevent overfitting to the LRLs. Figure 1 illustrates the effectiveness of Cooldown, which significantly accelerates convergence on the LRL due to a high temperature (aggressive upsampling of LRLs) at the beginning of training and reduces overfitting to the LRL due to the lowering the temperature during training.
@@
  • Motivated by our findings, we propose Cooldown, a method to adjust the sampling temperature during training on unbalanced datasets. We show the effectiveness of Cooldown  in multilingual settings.
@@ 2.1 Notations and Task Description
We consider a model trained on a collection of data $\mathcal{D}=\{x\}_{i=1}^{N}$ from $K$ domains $\mathcal{D}=\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup...\cup\mathcal{D}_{K}$. Here, “domain" refers to sources (Books, Wikipedia, code) for general language modeling or different languages (English, French, Swahili) in multilingual language modeling. The total training loss $\mathcal{J}(\mathcal{D})$ is the sum of the losses of each example $\mathcal{L}(x)$.
$\mathcal{J}(\mathcal{D})=\sum_{x\in\mathcal{D}}\mathcal{L}(x).$
@@
Naive aggregation often results in imbalanced performance across domains when high-resource domains dominate the aggregated loss. Scalarization solves this issue by assigning weights $\mathbf{w}=\{w_{i}\}_{i=1}^{K}$ to each domain and aggregates the weighted sum of individual losses:
$\mathcal{L}_{S}(\mathbf{w})=\mathbb{E}_{x\in\mathcal{D}}\left[w_{f(x)}\mathcal{L}(x)\right],$
@@
Instead of assigning weights to losses, we can also sample more frequently from the low-resource domain to achieve balanced training. Temperature sampling achieves this by adjusting the probabilities of selecting instances from different domains based on their sizes. The sampling probability vector $p$ of each domain is given by:
$\forall i\in\{1,2,...,K\}:\;p(i;\tau)=\frac{|\mathcal{D}_{i}|^{\frac{1}{\tau}}}{\sum_{j=1}^{K}|\mathcal{D}_{j}|^{\frac{1}{\tau}}},$
@@
$\mathcal{L}_{TS}(\tau)=\mathop{\mathbb{E}}_{\begin{subarray}{c}k\sim p(\cdot;\tau)\\ x\sim\mathcal{D}_{k}\end{subarray}}\Big{[}\mathcal{L}(x)\Big{]}$
The common understanding is that Temperature Sampling is mathematically equivalent to Scalarization (Choi et al., 2023; Xin et al., 2022), and we can use them interchangeably (Choi et al., 2023). In the next subsection, we will formalize this statement and show that these two are mathematically equivalent in full gradient descent. We will then show that they are not equivalent under stochastic gradient descent.
@@ 3.1 Theoretical Analysis
For any sampling temperature $\tau$, there exists a set of weights $\mathbf{w}_{\tau}=\{w_{1},w_{2},...,w_{K}\}$ for the Scalarization loss such that this loss is equivalent to the Temperature Sampling loss, both computed based on the whole data $\mathcal{D}$.
@@
$w_{i}=\frac{\sum_{j=1}^{K}|\mathcal{D}_{j}|}{|\mathcal{D}_{i}|}\cdot\frac{|\mathcal{D}_{i}|^{\frac{1}{\tau}}}{\sum_{j=1}^{K}|\mathcal{D}_{j}|^{\frac{1}{\tau}}}=\frac{p(i;\tau)}{p(i;1)},$
$\mathcal{L}_{S}(\mathbf{w}_{\tau})$ $=\mathop{\mathbb{E}_{x\sim\mathcal{D}}}\left[w_{f(x)}\mathcal{L}(x)\right]$
$=\sum_{x\in\mathcal{D}}\frac{p(f(x);1)}{|\mathcal{D}_{f(x)}|}w_{f(x)}\mathcal{L}(x)$
$=\sum_{i=1}^{K}p(i;\tau)\sum_{x\in\mathcal{D}_{i}}\frac{1}{|\mathcal{D}_{i}|}\mathcal{L}(x)$
$=\mathop{\mathbb{E}}_{i\sim p(\cdot;\tau)}\left[\sum_{x\in\mathcal{D}_{i}}\frac{1}{|\mathcal{D}_{i}|}\mathcal{L}(x)\right]$
$=\mathbb{E}_{i\sim p(\cdot;\tau)}[\mathbb{E}_{x\sim\mathcal{D}_{i}}[\mathcal{L}(x)]]=\mathcal{L}_{TS}(\tau).$
@@
For any $\tau$, let $\mathbf{w}_{\tau}$ be the set of weights such that $\mathcal{L}_{TS}(\tau)=\mathcal{L}_{S}(\mathbf{w}_{\tau})$. Let $\nabla\mathcal{L}(x)$ be the gradient with respect to a single datapoint $x$. We denote the stochastic gradient under Scalarization as $\nabla\mathcal{L}_{S}(x;\mathbf{w}_{\tau})=\{\nabla w_{f(x)}\mathcal{L}(x)|x\sim\mathcal{D}\}$, which is the gradient $\nabla w_{f(x)}\mathcal{L}(x)$ when a single sample $x$ is uniformly drawn from the dataset $\mathcal{D}$. Similarly, we denote the stochastic gradient under Temperature Sampling as $\nabla\mathcal{L}_{TS}(x;\tau)=\{\nabla\mathcal{L}(x)|x\sim\mathcal{D}_{i},i\sim p(i;\tau)\}$. Then both $\nabla\mathcal{L}_{S}(x;\mathbf{w}_{\tau})$ and $\nabla\mathcal{L}_{TS}(x;\tau)$ are unbiased estimates of the total gradient.
@@
By definition, we have
$\mathbb{E}[\nabla\mathcal{L}_{s}(x;\mathbf{w}_{\tau})]$ $=\mathop{\mathbb{E}}_{x\sim\mathcal{D}}\left[w_{f(x)}\nabla\mathcal{L}(x)\right]$
$=\mathbb{E}_{\mathcal{D}_{i}\sim p(\cdot;\tau)}[\mathbb{E}_{x\sim\mathcal{D}_{i}}[\nabla\mathcal{L}(x)]]$
$=\mathbb{E}[\nabla\mathcal{L}_{TS}(x;\tau)]$
@@
Using the same notation in Corollary 1.1, we have $\mathrm{Var}(\nabla\mathcal{L}_{S}(x;\mathbf{w}_{\tau}))\geq\mathrm{Var}(\nabla\mathcal{L}_{TS}(x;\tau))$.
We defer the proof of Theorem 2 to Appendix B.
@@
The difference
$\Delta=\mathrm{Var}(\nabla\mathcal{L}_{S}(x;\mathbf{w}_{\tau}))-\mathrm{Var}(\nabla\mathcal{L}_{TS}(x;\tau))$
monotonically increases when $\tau\geq 1$.
@@
Figure 2: Variance of Scalarization $\sum_{i}\frac{p(i;\tau)^{2}}{p(i;1)}$ by sampling temperature $\tau$. A large temperature or a skewed distribution of $\mathcal{D}$ induces a much larger variance for Scalarization. Distributions $\mathcal{D}_{i}\propto\frac{1}{i^{\alpha}}$. See Appendix C for details of the experiment setup.
@@ 3.2 Empirical Evidence
Figure 3: The distribution of gradient norm between mini-batches on En-{Cs, Ro} for Temperature Sampling and Scalarization. Scalarization induces a larger variance (2.25 >>> 0.62) between mini-batch gradient norms compared to Temperature Sampling, as indicated by Theorem 3.
@@
Empirical Validation of Theorem 2
We first validate that Temperature Sampling has a lower variance in gradient estimation than scalarization, as predicted by Theorem 2. Figure 3 illustrates the variance between mini-batch gradients for TS and scalarization, confirming that TS reduces gradient variance ($2.25\rightarrow 0.62$).
Next, we observe that this lower variance leads to faster convergence during training. Figure 4 shows the validation loss curves over training iterations, where TS consistently converges faster than scalarization across all temperatures we experimented with. Notably, the larger the temperature, the greater the gap in convergence speed between TS and Scalarization. This suggests that when there is a significant mismatch in data sizes, using Temperature Sampling with a high temperature is beneficial. Intuitively, low-resource languages (LRLs) with very little data should be upsampled more aggressively to accelerate convergence.
We summarize our findings below: